4. An antelope runs with a constant velocity of 22.0 ms- by a cheetah that is lying in the grass. The cheetah gets up and begins chasing directly after the antelope just as it has gone past, with a constant acceleration of 8.60 ms2, until it reaches its top speed of 30.0 ms- At tat point the cheetah runs with constant velocity. How far does the cheetah run before it catches up to the antelope?
Answer #1

Solution) initial speed u = 22m/s

Final speed v = 30m/s

Acceleration a = 8.60m/s^2

Distance S = ?

We have v^2 - u^2 = 2aS

(30)^2 - (22)^2 = 2(8.60)S

S = 24.18m